# A mechanism-informed reinforcement learning framework for shape optimization of airfoils

### v1

@@ 7 Conclusion
In this work, we present a novel framework that integrates reinforcement learning with goal-oriented PDEs-based systems for airfoil shape optimization. Operating within the challenging environment of unstructured meshes, we encountered significant obstacles such as mesh tangling, the management of high-dimensional state representations, and the occurrence of non-smooth geometries. To address these issues, we developed a suite of mesh recovery techniques aimed at improving mesh quality.
It is important to highlight that the reinforcement learning approach was meticulously tailored to align with the intrinsic motivation behind the optimization objectives. The design of the neural network was carefully crafted to meet these specific requirements while circumventing the potential for gradient vanishing issues. Moreover, we established a nuanced system of rewards and penalties to facilitate the interactive optimization process. Upon conducting experimental validations of shape optimization, our findings demonstrate that the proposed framework is adept at managing simulations with precisely calculated objective functions.
However, due to the current theoretical limitations of DWR-based mesh adaptation methods, more complex target functionals have not yet been further explored. Looking forward, we aim to integrate advanced techniques, such as the multi-mesh method for tackling more intricate target functionals and employing multi-agent strategies for enhanced exploration efficiency. We believe that multi-agent systems can achieve a deeper integration with traditional optimization algorithms based on adjoint equations, leading to more efficient exploration patterns. Furthermore, the adoption of various acceleration techniques, including the incorporation of additional learning modules, parallelization strategies, and more efficient network architectures, represents a promising direction for future development.

### v2

@@ 6 Conclusion
In this work, we present a mechanism-driven reinforcement learning framework. It integrates reinforcement learning with goal-oriented PDEs-based systems for shape optimization of airfoil. We developed a suite of mesh recovery techniques aimed at improving mesh quality to overcome the challenge brought by high-dimensional states and unstructured meshes. It is important to highlight that the reinforcement learning approach was meticulously tailored to align with the intrinsic motivation behind the optimization objectives. Upon conducting experimental validations of shape optimization, our findings demonstrate that the proposed framework is adept at managing simulations with precisely calculated objective functions.
However, more complex target functionals have not yet been further explored. Looking forward, we aim to integrate advanced techniques, such as the multi-mesh method for tackling more intricate target functionals and employing multi-agent strategies for enhanced exploration efficiency. It should be noted that multi-agent systems can achieve a deeper integration with traditional optimization algorithms based on adjoint equations, leading to more efficient exploration patterns. Furthermore, the adoption of various acceleration techniques, including the incorporation of additional learning modules, parallelization strategies, and more efficient network architectures, represents a promising direction for future development.
Acknowledgments
The authors would like to thank the support from The Science and Technology Development Fund, Macao SAR (No. 0082/2020/A2), National Natural Science Foundation of China (Nos. 11922120, 11871489), MYRG of University of Macau (No. MYRG-GRG2023-00157-FST-UMDF).